# Failure Modes of LLMs for Causal Reasoning on Narratives

### v2

@@ 3.3 Parametric Knowledge Conflict
We also find that LLMs tend to rely on parametric knowledge when it is present, and can fail when narratives run counter to their parametric knowledge. To test this, instead of picking random events, we have to draw out the LLM’s pre-existing parametric knowledge. To do so, we prompt the LLM to pick a series of events such that each event has some relation to the subsequent event–either the event is Causal to the next event (e.g., disease causes shorter lives) or the event is Anti-Causal (e.g., disease causes longer lives). Let’s call this graph of parametric knowledge $P$. We then take the odd indexed events (1st, 3rd etc) from $P$ and place them in the first half of the causal ground truth graph $G$ and the even indexed events (2nd, 4th etc) from $P$ in the second half of $G$. As we can see in the example illustration in Figure 3, each node in the first half of the causal graph $G$ now has a parametric relation to a specific node in the second half of $G$. For example we might know that node $1$ is Anti-Causal to node $2$ from parametric knowledge. Thus when we make the causal ground truth graph ($1\rightarrow 3\rightarrow$ 2), create a narrative from it, and then ask the LLM if node $1$ causes $2$ based on the narrative, it should say yes even though that disagrees with its parametric knowledge. After the ground truth graph is created, we generate the narrative in the Forward topological orientation to avoid confounding failure modes.
We find that, in synthetic experiments, this kind of occurrence is generally rare and the LLM finds the correct causal relation generally only when that relation agrees with its parametric knowledge. This is exemplified in the plot in Figure 3 where we observe that when certain pairs of events are Causal in the parametric knowledge, the LLM also generally finds that the pairs of events in question are Causal given the narrative. However, when the pairs of events are Anti-Causal (like nodes $1$ and $2$ in our previous example), even though the LLM should find them to be Causal given the narrative it generally does not.
@@
We also notice an interesting phenomenon for the Anti-Causal parametric case where using just the extracted graph provides massive improvements over any prompting strategy that involves using the narrative to directly answer questions. This strategy is comparable in performance to when the parametric knowledge is Causal. The extracted graph for the Claude 3.5 Sonnet experiment in the right side of 4 mirrors the same story. It seems that the narrative may only serve to distract the LLM when parametric knowledge disagrees with the narrative. When the narrative is provided, the LLM will often say things (in its COT explanation) like “Although this relationship is displayed in the chain of events in the narrative, it is not logical and counter-intuitive” and answer incorrectly. This happens despite explicit instruction to ignore outside information, allow for ill-logical relationships, or answer solely based on the hypothetical narrative.
3.4 Failure of Long-Term Causal Reasoning
@@
In conditions where the LLM exhibits failure modes (Reverse and Anti-Causal orientations), the performance also tends to decay as the size of the narrative and the number of events in the narrative increases. As we can see in Figures 2 and 3, it seems that the longer the narrative is, the more the LLM relies on shortcuts instead of performing reasoning. This finding is supported by the well known phenomenon that LLMs often fail to reason about longer form contexts especially as they “get lost in the middle” (Liu et al., 2023). What perhaps is a surprising result that is the extracted graph can often maintain a relatively consistent level of accuracy for causal answering across narrative sizes even when only using the narrative in the prompt leads to severe dips in accuracy.
@@ 4 Experiments with Real World Causal Graphs
In this section, we test our methods on narratives constructed using real-world causal graphs from CauseNet (Heindorf et al., 2020), a large-scale knowledge graph of (claimed) causal relationships between real-world concepts. We use GPT-4o (OpenAI et al., 2024) for our experiments.
4.1 Generating the narratives
The CauseNet dataset can be represented as a collection of $D$ tuples $\{(C_{i},E_{i},\mathbf{S}_{i}\}_{i=1}^{D}$, where $C_{i}$ denotes the cause (e.g., fatigue), $E_{i}$ denotes the effect (e.g., accidents), and $\mathbf{S}_{i}$ is a set of sentences (extracted from Wikipedia and ClueWeb12 (Callan, 2012)) that entail a causal relationship from $C_{i}$ to $E_{i}$. We retrieve causal chain graphs $V_{1}\rightarrow V_{2}\rightarrow\ldots\rightarrow V_{N}$ of various lengths, where each causal relation $V_{i}\rightarrow V_{i+1}$ is from CauseNet. Below, we describe the strategy for generating narratives for a given chain graph (see Appendix B.1 for the prompt templates).
@@ 4.1 Generating the narratives
Each chain graph is verbalized into a narrative by prompting the LLM to generate a sentence one edge at a time such that the sentence for $V_{i}\rightarrow V_{i+1}$ follows the sentence for the previously verbalized edge, ensuring that the narrative as a whole is sensible. We generate the narratives by enumerating the edges in the topological order of the graph and in the reverse order. We call this semi-synthetic because the nodes in the graph represent real-world causal relationships but the narratives are synthetic. For example, the following is the generated narrative in the forward direction for the causal chain fatigue $\rightarrow$ accidents $\rightarrow$ injury:
@@
Additional examples of semi-synthetic and real-world narratives are presented in Appendix B.2 (the entire set of narratives used for our experiments is available in the linked code).
4.2 Effect of Narrative Topological Ordering and Chain Length
As described in the previous section, we verbalize each causal chain graph $V_{1}\rightarrow V_{2}\rightarrow\ldots\rightarrow V_{N}$ from CauseNet into a narrative in the forward and reverse topological order. We give the narrative as input to the LLM and evaluate its causal reasoning abilities using the following prompting strategies (see Appendix B.3 for the prompt templates): (1) (Direct prompting) For every pair of nodes $(V_{i},V_{j})$, we ask does $V_{i}$ have a direct or indirect causal effect on $V_{j}$; (2) (CoT prompting) Same as (1), but we now use CoT and ask the LLM to think step-by-step; and (3) (Graph Extraction) Generate a causal chain graph faithful to the narrative when given the node identities $(V_{1},V_{2},\ldots,V_{N})$ in a random order.
We evaluate the accuracy for each pair of nodes $(V_{i},V_{j})$ for the three prompting strategies on the semi-synthetic and real-world narratives (see Fig. 5). We denote the Direct prompting strategy (1) using the labels Forward and Reverse; the CoT prompting strategy (2) using the labels Forward_COT and Reverse_COT; and the accuracy using the extracted graph from strategy (3) using the labels Forward_Graph and Reverse_Graph.
(a) Semi-synthetic narratives
(b) Real-world narratives
@@ 4.2 Effect of Narrative Topological Ordering and Chain Length
In both the semi-synthetic (Fig. 5(a)) and real-world narratives (Fig. 5(b)), the Forward_Graph strategy performs the best, with its accuracy remaining stable even as the chain length increases. By contrast, Reverse_Graph has relatively lower accuracy, with the accuracy declining as the chain length increases (this effect is particularly significant in the real-world narratives). We also observe that Forward(_COT) outperforms Reverse(_COT), with the Reverse(_COT) accuracy declining substantially as the chain size gets large. Additionally, unlike the forward direction, the graph extraction does not improve accuracy substantially over the Direct and CoT-based prompting strategies. Altogether, these results demonstrate that LLM’s ability to extract the causal graph as well as determine causality between two nodes declines when the narrative is verbalized in the reverse order, with the impact of the increasing chain lengths being particularly pronounced in the reverse direction.
4.3 Effect of Parametric Knowledge Conflicts.
Forward (w CoT) Reverse (w CoT) Forward Graph Reverse Graph
@@
Without Conflict 97.86 (97.94) 95.85 (96.47) 98.54 87.99
With Conflict 65.47 (80.54) 45.63 (67.83) 97.70 84.32
@@
Without Conflict 95.54 (96.06) 92.38 (91.63) 87.74 57.93
With Conflict 48.03 (59.57) 27.40 (49.43) 85.67 53.09
@@
Table 1: The average accuracy across different narratives with the three prompting strategies partitioned by whether the cause-effect pairs conflict with the LLM’s parametric knowledge (we omit the 95% CIs as they are smaller than $0.3$). The accuracy (with and without CoT) is substantially lower when knowledge conflicts exist, suggesting that LLM relies on this knowledge rather than the narrative. By contrast, the graph extraction is more robust to knowledge conflicts.
Next, we analyze the extent to which the LLM relies on its parametric knowledge to answer the causal query as opposed to the specific causal effects expressed in the narrative. For every pair of nodes $(V_{i},V_{j})$ in the chain graphs, we extract the parametric knowledge of the LLM by prompting the LLM to answer “Does $V_{i}$ typically have a causal effect on $V_{j}$?”. This parametric knowledge represents some average case notion of causality learned by the LLM from the pretraining corpora.
Successful causal reasoning often requires going beyond the parametric causal knowledge as it can differ from the causality between events in the context of the narrative at hand. For example, in a chain graph from the our dataset, there is a causal path from streambank erosion to higher prices, but this contradicts the LLM’s parametric knowledge since this causal effect may not typically exist in the real-world. In the chain graphs we evaluate, roughly $39\%$ of cause-and-effect pairs conflict with the LLM’s parametric knowledge.
We evaluate the three prompting strategies separately on the subsets of cause-and-effect pairs that are in agreement and in conflict with the parametric knowledge (see Table 1). We observe that when there is no conflict (i.e., the parametric knowledge agrees with the causality expressed in the narrative), the accuracies in both the Forward and Reverse directions (with and without CoT) are greater than $90\%$. However, when the parametric knowledge conflicts with the narrative’s causality, the accuracy of Forward and Reverse is significantly lower: even with CoT, the LLM has below-random accuracy in the reverse direction. This suggests that when asked to reason about cause and effect in a narrative, the LLM seems to rely heavily on its parametric knowledge and is unable to grasp the specific causal chains expressed in the narrative itself (despite the causal chains as a whole being realistic).
Interestingly, when using extracted graph for performing causal reasoning (Forward_Graph and Reverse_Graph), the performance in both cases seems to be comparable. This is likely because when asked to extract the graph from the narrative with the given set of nodes, the LLM pays more attention to the entire narrative as opposed to when directly queried on a cause-and-effect pair (where the LLM defaults to its parametric knowledge). These results show that even when the LLM constructs a reasonably good causal chain graph in the Forward direction, the LLM does not leverage this graph when queried directly about the causal effects in the narrative (even with CoT), further highlighting the advantage of extracting the causal graph directly.
Finally, we also observe that the accuracy of Forward and Reverse prompting strategies in Fig. 5 were largely driven by the parametric knowledge. To disentangle this effect, we evaluate the three prompting strategies across different chain lengths only on the subset of cause-and-effect pairs where a knowledge conflict exists (see Fig. 6). We observe that the accuracy of all prompting strategies, except Forward_Graph and Reverse_Graph, decline substantially relative to Fig. 5, with Reverse falling well below the random baseline of $50\%$ as the chain length gets larger. In the real-world narratives (Fig. 6(b)), even with CoT, the accuracy does not improve over the random baseline of $50\%$ in the reverse direction, demonstrating that the LLM ignores the causal chain in the narrative. Qualitatively, we observe that in the case of knowledge conflicts, the CoT explanation justifies the incorrect (according to the narrative) causal relationship by injecting its parametric knowledge into the chain-of-thought. Additionally, we also evaluate the three prompting strategies on the subset of cause-and-effect pairs in agreement with parametric knowledge (see Appendix B.4), showing high accuracy of the Direct and CoT-based prompting strategies on this subset.
(a) Semi-synthetic narratives
(b) Real-world narratives
Figure 6: Results on the subset of cause-and-effect pairs that conflict with the LLM’s parametric causal knowledge. We observe that while the graph extraction is robust to conflicts, the Direct and CoT prompting strategies suffer significantly (particularly in the reverse direction), indicating that the LLM returns its parametric knowledge instead of paying attention to the narrative.
@@ 5 Discussion
Our work takes some initial strides towards examining the reasons behind the success and failure of LLMs to reason causally on narratives that express a chain of causal events. We focus on two questions of key importance in causality: (1) Does one event cause another? and (2) Can the LLM extract the causal graph from the narrative. We find three significant failure modes of LLM reasoning by conducting experiments in carefully controlled synthetic, semi-synthetic and real-world settings:
  1. 1.
    Topological Ordering: LLMs tend to perform well when the ordering of events in the narratives matches that of the ordering of the underlying graph. For example if B is mentioned first in the narrative and then A, then the LLM would would expect $B\rightarrow A$ even if $A\rightarrow B$ is the truly underlying causal graph. As such, performance breaks down when the ordering of events in the narratives does not match that of the underlying graph.
  2. 2.
    Parametric Knowledge: LLMs rely on their parametric pre-training knowledge as a shortcut to infer causal relations. When the narrative suggests one causal relation (e.g., $A\rightarrow B$) and the parametric knowledge suggests another ($B\rightarrow A$), the LLM will often answer using its parametric knowledge, ignoring the narrative.
  3. 3.
    Narrative Size: LLM accuracy degrades as the chain length increasing, resulting in poor reasoning on longer narratives.
We also find that asking the LLM to extract what it predicts to be the underlying graph from the narrative, and using just that extracted graph or the extracted graph in concert with the narrative to answer can lead to mitigation of these failure modes.
Our work builds upon past work that examines causal reasoning in LLMs. Various works including Jin et al. (2023) suggest that LLMs may have difficulty performing causal reasoning in tasks involving numerical probabilistic computations. Previous work by Zečević et al. (2023) conjectured that LLMs simply imitate causal relations from their training data. Our work aims to assess causal reasoning from textual narratives (not just on individual sentences) that is required to understand why certain events occurred and how they would impacted by certain interventions. Like previous works, we find a strong reliance on pre-training knowledge. However, we also suggest a way to mitigate this by using the LLM to extract the entire causal graph from the narrative at once. When doing this, the graph remains generally accurate even in the presence of conflicts with parametric pre-training knowledge.

### v3

@@ 3.3 Impact of Parametric Knowledge (In)consistency
We also find that LLMs tend to rely on parametric knowledge when it is present, and can fail when narratives are inconsistent with the LLM’s parametric knowledge. To test this, we elicit the LLM’s pre-existing parametric knowledge when generating the event chains. We prompt the LLM to pick a series of events such that each event has some relation to the subsequent event – either the event is Causal to the next event (e.g., disease causes shorter lives) or the event is Anti-Causal (e.g., disease causes longer lives). For example, we might know that node $1$ is Anti-Causal to node $2$ from parametric knowledge. Thus, when we make the causal ground truth graph $1\rightarrow 3\rightarrow$ 2 (this disagrees with parametric knowledge), create a narrative from it, and then ask the LLM if node $1$ causes $2$ based on the narrative: it should say yes based on the narrative even though that disagrees with its parametric knowledge. After the ground truth graph is created, we generate the narrative in the Forward topological orientation to avoid confounding failure modes. The full process (along with illustration) explaining how the parametric and causal graphs are created is in Appendix A.2. As a textual example, assume that we know a parametric anti-causal link exists from stressful job to increased happiness, and from lack of sleep to improved cognitive function. We can then construct the causal chain Stressful JobLack of SleepIncreased HappinessImproved Cognitive Function. From this causal chain, we create the narrative:
The constant demands of a stressful job led to her experiencing chronic lack of sleep. Surprisingly, she found that the lack of sleep heightened her sense of euphoria, making her unusually cheerful at work. Increased happiness from this unexpected cheerfulness seemed to improve her cognitive function.
If the LLM is asked if a stressful job leads to increased happiness, the parametric knowledge shortcut indicates the answer should be no – however, the shortcut fails as the narrative indicates that a (indirect) causal link does exist.
Models Exploit Parametric Knowledge
@@
We find that, in synthetic experiments, the LLM finds the correct causal relation generally only when that relation agrees with its parametric knowledge. This is exemplified in the plot in Figure 2 (left) where we see good performance on narratives that agree with parametric knowledge (Causal parametric knowledge) and poor performance on narratives that disagree with parametric knowledge (Anti-Causal parametric knowledge). We also notice an interesting phenomenon for the Anti-Causal parametric case where using just the extracted graph provides massive improvements over any prompting strategy that involves using the narrative to directly answer questions. This strategy is comparable in performance to when the parametric knowledge is Causal. It seems that the narrative may only serve to distract the LLM when parametric knowledge disagrees with the narrative.
3.4 Impact of Narrative Complexity
In the previous sections, we identified two shortcuts which models exploit in causal reasoning tasks. Here, we test the influence of narrative complexity on these failure modes. We examine two measures of complexity: (a) the narrative length and (b) the presence of complex graph structures.
Narrative Length
@@ 3.4 Impact of Narrative Complexity
In conditions where the LLM exhibits failure modes (Reverse and Anti-Causal orientations), the performance also tends to decay as the size of the narrative and the number of events in the narrative increases. As we can see in Figures 1 and 2 (Left), it seems that the longer the narrative is, the more the LLM relies on shortcuts instead of performing reasoning. However, the extracted graph $G^{\prime}$ can often maintain a consistently high level of accuracy across narrative sizes even for cases when a failure mode would normally be exhibited.
Causal Graph Complexity
As the bulk of our work has focused on detecting the simplest failure modes possible, we studied narratives with an underlying chain graph structure. However, the presence of more complex causal structures in the narrative could exacerbate the existing failure modes or trigger novel failures. To study this, we create causal graphs utilizing two common causal structures: Forks (one node has a causal relationship to multiple other nodes) and Colliders (multiple nodes have a causal relationship to the same node). We generate narratives (the complete algorithm is described in Appendix A.3) such that each underlying causal graph contains at least one of these structures, and may randomly contain multiple such structures based on the size of the narrative. An example is shown in Figure 3.
Heavy RainfallPower OutageFlooded StreetsTraffic Jam
Narrative: The heavy rainfall not only caused a power outage in several neighborhoods but also led to flooded streets. The aftermath of the power outage (disabling traffic lights) and the flooded roads (blocking street access) caused a traffic jam.
Figure 3: Causal graph with story showing a fork (first sentence) and a collider (second sentence).
As can be seen in Figure 2 (right side), we find that while the LLM generally performs worse at reasoning about the complex narratives than simple narratives (with underlying chain graphs), the gap is very starkly less than can be seen in the other failure modes. This finding can be supported by (Dettki et al., 2025) which finds that GPT-4o reasons similarly to humans on a single sentence that describes one collider relation. Our work extends their work by using a long-form narrative based on a causal graph with potentially multiple colliders and forks instead of only one collider.
@@ 4 Experiments with Real World Causal Graphs
In this section, we extend our analysis to narratives involving real-world causal graphs from CauseNet (Heindorf et al., 2020), a large-scale knowledge graph of (claimed) causal relationships between real-world concepts. We perform experiments using the GPT-4o (OpenAI et al., 2024) and Llama-3.1 8B models for our experiments. We concentrate our analysis on the same factors (positional biases and parametric knowledge consistency) as explored in the semi-synthetic settings.
4.1 Experimental Setting
The CauseNet dataset can be represented as a collection of $D$ tuples $\{(C_{i},E_{i},\mathbf{S}_{i}\}_{i=1}^{D}$, where $C_{i}$ denotes the cause (e.g., fatigue), $E_{i}$ denotes the effect (e.g., accidents), and $\mathbf{S}_{i}$ is a set of sentences (extracted from Wikipedia and ClueWeb12 (Callan, 2012)) that entail a causal relationship from $C_{i}$ to $E_{i}$. We retrieve causal chain graphs $V_{1}\rightarrow V_{2}\rightarrow\ldots\rightarrow V_{N}$ of various lengths, where each causal relation $V_{i}\rightarrow V_{i+1}$ is from CauseNet and verbalize these chains as narratives in the following ways:
@@ 4.1 Experimental Setting
In this setting, we use real causal graphs from CauseNet but synthetically verbalize them via the LLM. In particular, we prompt the LLM to generate sentences for each edge ($V_{i}\rightarrow V_{i+1}$) in the causal graph, while ensuring the sensibility of the entire narrative. For eammple, the following is a narrative for the chain fatigue $\rightarrow$ accidents $\rightarrow$ injury:
@@
Additional examples of semi-synthetic and real-world narratives are presented in Appendix C.1 (the entire set of narratives used for our experiments is available in the linked code).
Prompting Strategies
For simplicity, we limit the prompting techniques used to (see Appendix C.2 for the prompt templates): Standard QA Prompting, Chain-of-Thought and Explicit Causal Graph Extraction. We evaluate the accuracy for each pair of nodes $(V_{i},V_{j})$ for the three prompting strategies on the semi-synthetic and real-world narratives.
4.2 Impact of Event Ordering and Chain Length
As described in the previous section, we verbalize each causal chain graph $V_{1}\rightarrow V_{2}\rightarrow\ldots\rightarrow V_{N}$ from CauseNet into a narrative in the forward and reverse topological order.
@@ 4.2 Impact of Event Ordering and Chain Length
In both the semi-synthetic (Fig. 4 left) and real-world narratives (Fig. 4 right), the Forward Graph strategy performs the best, with its accuracy remaining stable even as the chain length increases. We observe that Forward Standard and CoT outperforms Reverse Standard and CoT, with the Reverse accuracy declining substantially as the chain size gets large. We also see that in this regime, extracting the causal graph makes inference in the Reverse orientation competitive with inference in Forward.
4.3 Effect of Parametric Knowledge Consistency
Experiment Setup
Next, we analyze the extent to which the LLM relies on its parametric knowledge to answer causal reasoning queries as opposed to the causal structure expressed in the narrative. For every pair of nodes $(V_{i},V_{j})$ in the chain graphs, we elicit the parametric knowledge of the LLM by asking the LLM whether a causal effect between the two nodes would be atypical (see Appendix B.2 for the exact prompts utilized). Through these prompts, we identify cause and effect chains which contradict the model’s parametric knowledge. For example, in a chain graph from our dataset, there is a path from streambank erosion to higher prices, but this contradicts the LLM’s parametric knowledge since this causal effect may not typically exist in the real-world. In total, we find that roughly 5 percent of the relations in CauseNet violate the LLM’s pretraining knowledge. We sampled narratives from CauseNet until we got 100 (of chain sizes between 3 and 9) narratives that contain relations that violate the LLM’s pre-training knowledge and 100 that are consistent. These narratives are constructed in the Forward topological ordering to avoid confounding failure modes.
LLM Performance Suffers on Atypical Causal Relations
We evaluate the three prompting strategies separately on the subsets of cause-and-effect pairs that are in agreement and in conflict with the parametric knowledge (see Table 1). We observe that when there is no conflict (i.e., the parametric knowledge agrees with the causality expressed in the narrative), the accuracies with and without CoT are greater than $90\%$. However, when the parametric knowledge conflicts with the narrative’s causality, the accuracy is significantly lower, even with CoT. This suggests that when asked to reason about cause and effect in a narrative, the LLM seems to rely heavily on its parametric knowledge and is unable to grasp the specific causal chains expressed in the narrative itself (despite the causal chains as a whole being realistic).
Explicit Causal Graph Extraction Avoids Shortcuts
Interestingly, when using extracted graph for performing causal reasoning, the performance is very high, both with and without conflicts. This is likely because when asked to extract the graph from the narrative, the LLM pays more attention to the entire narrative as opposed to when directly queried on a cause-and-effect pair (where the LLM defaults to its parametric knowledge). These results show that even when the LLM constructs a reasonably good causal chain graph, the LLM does not leverage this graph when queried directly about the causal effects in the narrative (even with CoT), further highlighting the advantage of extracting the causal graph directly.
Standard CoT Graph
@@ 4.3 Effect of Parametric Knowledge Consistency
Without Conflict 99.8 99.6 99.9
With Conflict 67.2 73.1 98.7
@@
Without Conflict 90.9 89.2 97.9
With Conflict 52.1 57.6 93.2
@@
Table 1: The average accuracy across different narratives with the three prompting strategies partitioned by whether the cause-effect pairs conflict with the LLM’s parametric knowledge (we omit the 95% CIs as they are smaller than $0.3$).
[…]
4.4 Narrative Complexity
We can see from Figure 4 that LLM performance degrades with narrative length, especially when a failure mode is present. We furthermore experimented with complex narratives with causal graphs containing forks and colliders (full graph and narrative creation algorithm in Appendix B.3). We can see in Figure 5, that in both the semi-synthetic and real-world settings that complex narratives (with colliders and forks) perform worse than simple narratives that have a causal chain graph as the ground truth. This gap ,while clear and noticeable, isn’t as stark as failure from parametric knowledge conflict (Table 1) or topological ordering (Figure 4). We do furthermore note that this is one area where extracting an explicit causal graph does not seem to significantly improve performance.
@@ 5 Discussion
Our work takes initial strides towards examining the success and failure of LLMs to reason causally on narratives that express causal events. We focus on two questions of key importance in causality: (1) Does one event cause another? (2) Can the LLM extract the causal graph from the narrative. We find three significant failure modes of LLM reasoning by conducting experiments in carefully controlled synthetic, semi-synthetic and real-world settings: Firstly, we find that LLMs rely heavily on topological ordering, performing well when the ordering of events in the narratives matches that of the ordering of the underlying causal graph. Secondly, we find that LLMs rely on their parametric knowledge as a shortcut to infer causal relations. Finally, we examine the role of causal structure complexity, finding that LLM accuracy degrades as the narrative length increases. Furthermore, LLMs perform slightly worse on reasoning when narratives contain structures such as colliders and forks. Beyond these failure modes, we show that more reliable causal reasoning can be elicited by prompting the LLM to explicitly identify the causal graph.